# MDBench: Benchmarking Data-Driven Methods for Model Discovery

### v2

@@ B.1 Advection
The advection system, adopted from (takamoto2022pdebench), is a simple linear system in 1D space with the equation
@@ B.2 Burgers
The Burgers equation, from (takamoto2022pdebench), models the diffusion process in fluid dynamics as
@@ B.3 Korteweg-de Vries (KdV)
The KdV system, taken from (brunton2016discovering), models unidirectional propagation of small-amplitude waves in shallow water. The governing equation is given by
@@
B.4 Kuramoto-Sivishinky (KS)
The KS equation is known for modeling the dynamics of various physical systems such as fluctuations in fluid films on inclines (brunton2016discovering). This equation includes a fourth-order diffusion term as a diffusive regularization of non-linear dynamics. The system is governed by the equation
@@ B.6 Reaction-Diffusion
Reaction diffusion systems are used extensively to study pattern forming systems in physics (brunton2016discovering). One particular class of DR systems with nonlinear coupling terms are governed by the coupled PDE
@@ B.7 Advection-Diffusion
The Advection-Diffusion system, adopted from (both2021deepmod), appears in heat transfer, pollution transport, and chemical physics, to name a few. We study the 2D Advection-Diffusion system
@@ B.8 Heat (Solar)
The Heat (Solar) dataset from (logg2012automated) simulates the heating of the earth’s surface by solar radiation. For a given dimension $d\in\{1,2,3\}$, a rectangular domain $\omega$ is defined with $x_{d-1}=0$ representing the earth’s surface. Using default settings, the domain will be 1.5m in depth and 0.75m in width (for 2D and 3D problems). In the 3D case, we sample the first 20 time points.
@@ B.9 Heat (Laser)
The Heat (Laser) dataset from (abali2016computational) simulates the dynamics of heat transfer through a rigid body with a custom heat source. In this dataset, the heat supply comes from a moving laser beam that is used to heat a 3D plate in a concentrated manner. The power of the laser follows
@@ B.10 Navier-Stokes Channel (NS Channel)
The incompressible Navier-Stokes equations model the motion of viscous fluids. The NS Channel dataset, adopted from (langtangen2017solving), simulates a liquid traveling through a 2D channel along the $x$ direction. The dynamics follow the system of equations
@@ B.11 Navier-Stokes Cylinder (NS Cylinder)
The NS Cylinder dataset is another example of a Navier-Stokes system as published in the FEATFLOW Benchmark Suite (featflow). The system records the velocity and pressure generated by flow through a 2D rectangular channel 2.2 m in length and 0.41 m in width, with a 5 cm wide cylinder obstruction centered at $xy$-coordinates $(0.2,0.2)$. In this system, the fluid density $\rho=1$ and the kinematic velocity $\mu=0.001$. With a maximum velocity of 1.5 m/s, the system has a Reynolds number $Re=100$, which results in turbulent flow, with vortex shedding beyond the cylinder.
@@
The unstructured mesh for the NS Cylinder dataset was generated using the mshr subpackage of FEniCS (fenics). The dataset simulates the system over 5 seconds, saving the data at 1,000 $Hz$ for a total of 5,000 timesteps. The dataset can be reproduced from the script fenics_navier_stokes_cylinder.py. Please note that the dataset must be converted into a structured mesh on a uniform grid prior to use for model discovery. The structured data can be generated from the unstructured version using the ns_cylinder_structured.py script. In our analysis, we subsample a rectangular grid in front of the cylinder (Figure A6) in space and $t\in[1,2],\Delta t=0.02$ in time.
@@ B.12 Reaction-Diffusion (Cylinder)
The Reaction-Diffusion Cylinder dataset utilizes the velocity component $\mathbf{v}$ of the solution to the NS Cylinder system to simulation an advection-diffusion-reaction system as detailed in (langtangen2017solving). Two chemical reagents $A$ and $B$ are carried by the fluid flow, i.e., advected, around the the cylinder and react to product chemical $C$, which is then disbursed downstream. The dynamical system of the chemical reactions is:
@@ Appendix C Methods and Hyperparameter Optimization
SINDy (brunton2016discovering) constructs a library of candidate functions (e.g., polynomials, trigonometric functions), evaluated on the state variables over time, and applies LASSO ($\ell_{1}$)-regularized regression to identify a sparse set of terms that best represent an ODE system. PDEFind (rudy2017data) extends SINDy to PDEs by evaluating the candidate library on a spatiotemporal grid.
WSINDy (messenger2021weak) is another linear method for PDE discovery that avoids explicit pointwise derivative estimation by leveraging the weak form of the governing equations. ESINDy and EWSINDy (fasel2022ensemble) extend SINDy and WSINDy, respectively, using bootstrap aggregation for improved robustness on ODE and PDE systems. All SINDy-based methods are implemented using the PySINDy library https://pysindy.readthedocs.io/en/latest/.
DeepMoD (both2021deepmod) is a linear approach that integrates a neural network to simultaneously learn the underlying function, its derivatives, and the coefficients of a predefined candidate library. Bayesian (more2023bayesian) builds on PDEFind by incorporating variational Bayes and a spike-and-slab prior for sparse model selection. Its implementation is available at https://github.com/TapasTripura/Bayesian-Discovery-of-PDEs.
PySR (cranmer2023interpretable) is a genetic programming method with an optimized Julia backend, enabling high-performance model discovery through a parallel population-based evolve-simplify-optimize loop. See the official repository at https://github.com/MilesCranmer/PySR. Operon (burlacu2020operon) is another efficient genetic programming method, implemented in C++ with vectorized operations and a compact linear tree representation. Its codebase is available at https://github.com/heal-research/operon.
EQL (martius2016extrapolation) is a deep learning-based method that embeds symbolic operations (e.g., addition, cosine) into the activation functions of a shallow neural network. After training, the network weights correspond to the coefficients of the discovered equation. We use the implementation available at https://github.com/cavalab/eql. uDSR (landajuela2022unified) provides a unified framework that integrates multiple approaches, from linear models to genetic programming, for model discovery. Its code is available at https://github.com/dso-org/deep-symbolic-optimization.
End2end (kamienny2022end) is a general-purpose model discovery approach based on a sequence-to-sequence Transformer architecture, pretrained on large datasets of synthetic equations. The implementation can be found at https://github.com/facebookresearch/symbolicregression. ODEFormer (d2024odeformer) extends End2end to ODE systems. Unlike most generic model discovery methods, ODEFormer does not rely on derivative estimates and instead takes state trajectories and timestamps as input. Its implementation is available at https://github.com/sdascoli/odeformer.
@@
basis functions
[polynomial], [polynomial, sin, cos]
[polynomial, sin, cos, exp]
@@ D.1 ODE Systems
Figure 10(a) presents aggregated performance of model discovery methods on all ODE systems, as measured by NMSE of predicted derivatives on the test set. While LM tend to discover less complex equations as the noise level increases, GP methods and ODEFormer exhibit the opposite trend. Figures A12 and A13 provide more detailed metrics on how different noise levels affect the quality of equations for each dataset. The dataset names are described in Table A4. For more details about the equations and initial values, refer to (d2024odeformer) or scripts/strogatz_ode.py in the MDBench project repository.

### v3

@@ B.1 Advection
The advection system, adopted from (Takamoto et al. 2022), is a simple linear system in 1D space with the equation
@@ B.2 Burgers
The Burgers equation, from (Takamoto et al. 2022), models the diffusion process in fluid dynamics as
@@ B.3 Korteweg-de Vries (KdV)
The KdV system, taken from (Brunton et al. 2016), models unidirectional propagation of small-amplitude waves in shallow water. The governing equation is given by
@@
B.4 Kuramoto-Sivashinsky (KS)
The KS equation is known for modeling the dynamics of various physical systems such as fluctuations in fluid films on inclines (Brunton et al. 2016). This equation includes a fourth-order diffusion term as a diffusive regularization of non-linear dynamics. The system is governed by the equation
@@ B.6 Reaction-Diffusion
Reaction diffusion systems are used extensively to study pattern forming systems in physics (Brunton et al. 2016). One particular class of DR systems with nonlinear coupling terms are governed by the coupled PDE
@@ B.7 Advection-Diffusion
The Advection-Diffusion system, adopted from (Both et al. 2021), appears in heat transfer, pollution transport, and chemical physics, to name a few. We study the 2D Advection-Diffusion system
@@ B.8 Heat (Solar)
The Heat (Solar) dataset from (Logg et al. 2012) simulates the heating of the earth’s surface by solar radiation. For a given dimension $d\in\{1,2,3\}$, a rectangular domain $\omega$ is defined with $x_{d-1}=0$ representing the earth’s surface. Using default settings, the domain will be 1.5m in depth and 0.75m in width (for 2D and 3D problems). In the 3D case, we sample the first 20 time points.
@@ B.9 Heat (Laser)
The Heat (Laser) dataset from (Abali 2016) simulates the dynamics of heat transfer through a rigid body with a custom heat source. In this dataset, the heat supply comes from a moving laser beam that is used to heat a 3D plate in a concentrated manner. The power of the laser follows
@@ B.10 Navier-Stokes Channel (NS Channel)
The incompressible Navier-Stokes equations model the motion of viscous fluids. The NS Channel dataset, adopted from (Langtangen and Logg 2017), simulates a liquid traveling through a 2D channel along the $x$ direction. The dynamics follow the system of equations
@@ B.11 Navier-Stokes Cylinder (NS Cylinder)
The NS Cylinder dataset is another example of a Navier-Stokes system as published in the FEATFLOW Benchmark Suite (Turek and Becker 1998). The system records the velocity and pressure generated by flow through a 2D rectangular channel 2.2 m in length and 0.41 m in width, with a 5 cm wide cylinder obstruction centered at $xy$-coordinates $(0.2,0.2)$. In this system, the fluid density $\rho=1$ and the kinematic velocity $\mu=0.001$. With a maximum velocity of 1.5 m/s, the system has a Reynolds number $Re=100$, which results in turbulent flow, with vortex shedding beyond the cylinder.
@@
The unstructured mesh for the NS Cylinder dataset was generated using the mshr subpackage of FEniCS (Alnæs et al. 2015). The dataset simulates the system over 5 seconds, saving the data at 1,000 $Hz$ for a total of 5,000 timesteps. The dataset can be reproduced from the script fenics_navier_stokes_cylinder.py. Please note that the dataset must be converted into a structured mesh on a uniform grid prior to use for model discovery. The structured data can be generated from the unstructured version using the ns_cylinder_structured.py script. In our analysis, we subsample a rectangular grid in front of the cylinder (Figure A6) in space and $t\in[1,2],\Delta t=0.02$ in time.
@@ B.12 Reaction-Diffusion (Cylinder)
The Reaction-Diffusion Cylinder dataset utilizes the velocity component $\mathbf{v}$ of the solution to the NS Cylinder system to simulation an advection-diffusion-reaction system as detailed in (Langtangen and Logg 2017). Two chemical reagents $A$ and $B$ are carried by the fluid flow, i.e., advected, around the the cylinder and react to product chemical $C$, which is then disbursed downstream. The dynamical system of the chemical reactions is:
@@ Appendix C Methods and Hyperparameter Optimization
SINDy (Brunton et al. 2016) constructs a library of candidate functions (e.g., polynomials, trigonometric functions), evaluated on the state variables over time, and applies LASSO ($\ell_{1}$)-regularized regression to identify a sparse set of terms that best represent an ODE system. PDEFind (Rudy et al. 2017) extends SINDy to PDEs by evaluating the candidate library on a spatiotemporal grid.
WSINDy (Messenger and Bortz 2021) is another linear method for PDE discovery that avoids explicit pointwise derivative estimation by leveraging the weak form of the governing equations. ESINDy and EWSINDy (Fasel et al. 2022) extend SINDy and WSINDy, respectively, using bootstrap aggregation for improved robustness on ODE and PDE systems. All SINDy-based methods are implemented using the PySINDy library https://pysindy.readthedocs.io/en/latest/.
DeepMoD (Both et al. 2021) is a linear approach that integrates a neural network to simultaneously learn the underlying function, its derivatives, and the coefficients of a predefined candidate library. Bayesian (More et al. 2023) builds on PDEFind by incorporating variational Bayes and a spike-and-slab prior for sparse model selection. Its implementation is available at https://github.com/TapasTripura/Bayesian-Discovery-of-PDEs.
PySR (Cranmer 2023) is a genetic programming method with an optimized Julia backend, enabling high-performance model discovery through a parallel population-based evolve-simplify-optimize loop. See the official repository at https://github.com/MilesCranmer/PySR. Operon (Burlacu et al. 2020) is another efficient genetic programming method, implemented in C++ with vectorized operations and a compact linear tree representation. Its codebase is available at https://github.com/heal-research/operon.
EQL (Martius and Lampert 2016) is a deep learning-based method that embeds symbolic operations (e.g., addition, cosine) into the activation functions of a shallow neural network. After training, the network weights correspond to the coefficients of the discovered equation. We use the implementation available at https://github.com/cavalab/eql. uDSR (Landajuela et al. 2022) provides a unified framework that integrates multiple approaches, from linear models to genetic programming, for model discovery. Its code is available at https://github.com/dso-org/deep-symbolic-optimization.
End2end (Kamienny et al. 2022) is a general-purpose model discovery approach based on a sequence-to-sequence Transformer architecture, pretrained on large datasets of synthetic equations. The implementation can be found at https://github.com/facebookresearch/symbolicregression. ODEFormer (d’Ascoli et al. 2024) extends End2end to ODE systems. Unlike most generic model discovery methods, ODEFormer does not rely on derivative estimates and instead takes state trajectories and timestamps as input. Its implementation is available at https://github.com/sdascoli/odeformer.
@@
basis functions [polynomial], [polynomial, sin, cos] [polynomial, sin, cos, exp]
@@ D.1 ODE Systems
Figure 10(a) presents aggregated performance of model discovery methods on all ODE systems, as measured by NMSE of predicted derivatives on the test set. While LM tend to discover less complex equations as the noise level increases, GP methods and ODEFormer exhibit the opposite trend. Figures A12 and A13 provide more detailed metrics on how different noise levels affect the quality of equations for each dataset. The dataset names are described in Table A4. For more details about the equations and initial values, refer to (d’Ascoli et al. 2024) or scripts/strogatz_ode.py in the MDBench project repository.